tickets at a local movie were sold at $6.00 for adults and $4.50
for students. if 264 tickets were sold for a total of $1563.00, how many adults tickets were sold ?

Answer by checkley75(3666)

A+S=264 OR S=264-A
6A+4.5S=1563.
NOW SUBSTITUTE (264-A) FOR S IN THIS EQUATION & SOLVE FOR A:
6A+4.5(264-A)=1563
6A+1188-4.5A=1563
1.5A=1563-1188
1.5A=375
A=375/1.5
A=250 NUMBER OF ADULT TICKETS SOLD.
264-250=14 NUMBER OF STUDENT TICKETS SOLD.
PROOF
6*250+4.5*14=1563
1500+63=1563
1563=1563